OpenStudy (anonymous):

Find the zeros of f(x)=2x^4-5x^3+11x^2-35x-21

OpenStudy (anonymous):

2x^4-5x^3+11x^2-35x-21 factor by grouping. (2x^4-5x^3)(11x^2-35x) - 21 x^3(2x-1) x(11x-35) - 21 (x^3+x- 21)(11x-35)(2x-1) factor again by grouping (x^3+x- 21) x(x^2+1)-21 x = 21 x^2 = -1 after squarerooting both sides you're left with imaginary numbers as well: x=+-1i (11x-35) 11x=35 x = (35/11) (2x-1) 2x=1 x=(1/2) so your zeros are: x=(1/2) x = (35/11) x=-1i x=+1i x = 21